### v1

@@ 4.4 System calibration
4.5 Single photon detection and coincidence counting
In the 93 km entanglement distribution experiment, the single photons are detected with SNSPDs, with the efficiency for idler channel at 45% while efficiency for signal channel is at 78%. Single photon count rate is 4.8 kcps for signal arm and 5.6 Mcps for idler arm. The delay cause by the deployed fiber is 457,369,970 ps which is equal to the distance measured by OTDR. The coincidence window used in local measurement is 200 ps while with deployed fiber is 60 ps for accidental suppression and higher CAR. The narrow does result in a lower photon pair rate. Reducing the background noise (detector dark counts, Raman noise in the source, and stray light coupling into the fiber network) should further improve the CAR and fidelity. When the deployed fiber distance is extended to 155 km, the singles count rate on signal and idler arm are around 15 kcps and 40 kcps respectively. Although this low single count rate will not saturate the detector, the SNSPD are set to 56% efficiency to reduce the dark count to 10 cps. The optical delay caused by the 62 km fiber is 302,113,173 ps which is equal to the length characterization by OTDR.
4.6 Fiber link information
A measurement of polarization stability of deployed fibers over a continuous period of 8 hours is conducted and the polarization state rotates (93 km link) and (62 km link) on the Poincare sphere. The results in Fig. 7. (a, b) indicate a relatively stable polarization behavior.
@@ 4.6 Fiber link information
Acknowledgements
This research is supported by the National Research Foundation, Singapore, and A*STAR under its CQT Bridging Grant and Quantum Engineering Programme (Award No. NRF2021-QEP2-01-P02), QEP2 grant: NRF2021-QEP2-01-P01, and US Air Force grant (FA2386-21-1-0076). D. Tan acknowledges funding from the National Research Foundation, Singapore, and A*STAR under its QEP 2.0 Programme (Award No. NRF2022-QEP2-01-P08), the National Research Foundation Investigatorship (NRF-NRFI08-2022-0003) and the Ministry of Education Tier 2 Grant (Award No. MOE2019-T2-2-178). We acknowledge funding support from the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme (National Quantum-Safe Network, NRF2021-QEP2-04-P01) and Netlink Trust for the provisioning of the fiber network. The authors acknowledge the assistance of setup alignment and literature review from Arya Chowdhury, and Lim En Teng, and consultation from Dr. Tanvirul Islam.
Declarations
The authors declare no conflicts of interest.

### v2

@@ METHODS
Single photon detection and coincidence counting
The entangled photons are detected by SNSPDs. The detection efficiency of the SNSPDs is tunable via the device bias current: higher bias increases both detection efficiency and dark count rate, but can reduce the maximum detectable count rate by making the detector more susceptible to latching.
For source characterization measurements, the SNSPD efficiency was intentionally set to a low value (45% and 49%) to increase the pump power substantially without saturating the detectors.
In the 93 km entanglement distribution experiment, the single photons were detected with SNSPDs set to higher efficiency: 78% for the signal channel and 45% for the idler channel. The single-photon count rates were 4.8 kcps for the signal arm and 5.6 Mcps for the idler arm. The delay caused by the deployed fiber is 457,369,970 ps, matching the distance measured by OTDR. The coincidence window used in local measurements is 200 ps, while with deployed fiber is 60 ps for accidental suppression and higher CAR. The narrow results in a lower photon-pair rate. Reducing the background noise (detector dark counts, Raman noise in the source, and stray light coupling into the fiber network) should further improve the CAR and fidelity.
When the deployed fiber distance was extended to 155 km, the singles count rates on the signal and idler arms were around 15 kcps and 40 kcps, respectively. Although this low single-photon count rate does not saturate the detector, the SNSPDs were set to 56% efficiency by lowering the bias current to reduce the dark count rate to 10 cps. The optical delay caused by the 62 km fiber is 302,113,173 ps, consistent with the length characterization by OTDR.
Fiber link information
A measurement of polarization stability of deployed fibers over a continuous period of 8 hours is conducted and the polarization state rotates (93 km link) and (62 km link) on the Poincaré sphere. The results in Fig. 7. (a, b) indicate a relatively stable polarization behavior.
@@
RESOURCE AVAILABILITY
Lead contact
Requests for further information and resources should be directed to and will be fulfilled by the lead contact Alexander Ling (alexander.ling@nus.edu.sg).
Materials Availability
This study did not generate new unique materials.
Data and code availability
The data reported in this study are available from the lead author upon reasonable request.
This paper does not report any original code.
Any additional information required to reanalyze the data reported in this paper is available from the lead contact upon request.
ACKNOWLEDGMENTS
This research is supported by the National Research Foundation, Singapore, and A*STAR under its CQT Bridging Grant and Quantum Engineering Programme (Award No. NRF2021-QEP2-01-P02), QEP2 grant: NRF2021-QEP2-01-P01, and US Air Force grant (FA2386-21-1-0076). D. Tan acknowledges funding from the National Research Foundation, Singapore, and A*STAR under its QEP 2.0 Programme (Award No. NRF2022-QEP2-01-P08), the National Research Foundation Investigatorship (NRF-NRFI08-2022-0003) and the Ministry of Education Tier 2 Grant (Award No. MOE2019-T2-2-178). We acknowledge funding support from the National Research Foundation, Singapore and A*STAR under its Quantum Engineering Programme (National Quantum-Safe Network, NRF2021-QEP2-04-P01) and Netlink Trust for the provisioning of the fiber network. The authors acknowledge the assistance of setup alignment and literature review from Arya Chowdhury, and En Teng Lim, and consultation from Dr. Tanvirul Islam, Dr. Haoran Zhang, and Dr. Shihao Ru.
AUTHOR CONTRIBUTIONS
J.D., X.Z., D.T.H.T., and A.L. conceived the experiment; G.F.R.C. and H.G. contributed to materials processing and fabrication; J.D. performed chip characterization; J.D. and X.Z. conducted entanglement measurements and analyzed the data; J.D., D.T.H.T., and A.L. wrote the manuscript; A.L. and D.T.H.T. supervised the project and acquired funding.
DECLARATION OF INTERESTS
Information related to the edge coupler design used in this work is covered by a pending international patent application (Application No. PCT/SG2024/050518, titled “Coupling Platform for Photonic Chips”) filed by Jinyi Du, Dawn T. H. Tan, and Alexander Ling.